Now I have an elimination problem.
x+2y=-7
-x+2y=1
I have tried setting it up like this: x+2y=-7
+ -x+2y= 1
----------- and I tried multipling equ 1
4y=-6 by 1 and equ 2 by -2

Answer by stanbon(48568)

I assume you made a typo on your 1st equation:
1) x+2y=-7
2)-x+2y=1
Add the two equations to get:
3) 4y=-6
y=-3/2
Substitute that value into 1) to get:
x+2(-3/2)=-7
x-3=-7
x=-4
Therefore the solution is (-4,-3/2)
Cheers,
stan H.